Compute the value of 9x2 + 4y2 if xy = 6 and 3x + 2y = 12.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Consider the equation 3x + 2y = 12

Now, square both sides:

(3x + 2y)^2= 122

=> 9x^2+ 12xy + 4y^2= 144

=>9x^2+ 4y^2= 144 – 12xy

since, xy = 6


9x^2+ 4y^2= 144 – 72

Thus, the value of 9x^2+ 4y^2= 72

Thanks